OpenStudy (anonymous):

if x^4 + 1/x^4 = 146 , thenwhich number is subtracyed from both sides to make LHS (x^2 + 1/x^2)^2 ?

ganeshie8 (ganeshie8):

expand (x^2 + 1/x^2)^2

OpenStudy (anonymous):

can you please give me a detailed solution?

ganeshie8 (ganeshie8):

do you knw (a+b)^2 formula, can you do what i asked above

OpenStudy (anonymous):

ya (a+b)^2=a^2+2ab+b^2

ganeshie8 (ganeshie8):

use that and expand (x^2 + 1/x^2)^2

OpenStudy (anonymous):

okk x^2+1/x^2)^2=x^4+2+1/x^4

ganeshie8 (ganeshie8):

like this :- \(\large (x^2 + \frac{1}{x^2})^2\) \(\large (x^2)^2 + (\frac{1}{x^2})^2 + 2.x^2.\frac{1}{x^2} \) \(\large x^4 + \frac{1}{x^4} + 2.\cancel{x^2}.\frac{1}{\cancel{x^2}} \) = ?

ganeshie8 (ganeshie8):

so we need to add what number ?

OpenStudy (anonymous):

-2

ganeshie8 (ganeshie8):

no, +2

OpenStudy (anonymous):

ok thanks

ganeshie8 (ganeshie8):

np, you're welcome :)

OpenStudy (anonymous):

but my qs is about subtracting can you please take a look at it?

OpenStudy (anonymous):

you there sir

ganeshie8 (ganeshie8):

then you're right !

OpenStudy (anonymous):

Sir can you please solve and show me entire thing because i am stll confused

ganeshie8 (ganeshie8):

i dont think so, maybe you need to go try multiple times until you feel comfortable

OpenStudy (anonymous):

can you help me with another question?

ganeshie8 (ganeshie8):

sure il be happy to help if i can :) ask..

OpenStudy (anonymous):

What height of cubical vessal if volume is 10 litres and base area is 1000 cm2

ganeshie8 (ganeshie8):

volume = base area * height 10 liter = 1000cm^2 * height 10 * 1000 cm^3 = 1000cm^2 * height

ganeshie8 (ganeshie8):

you can find the height in cm

OpenStudy (anonymous):

10 is the answer?

ganeshie8 (ganeshie8):

yes 10cm

ganeshie8 (ganeshie8):

let me knw if u dont get liter to cm^3 conversion..

OpenStudy (anonymous):

hws that?

ganeshie8 (ganeshie8):

hows what

OpenStudy (anonymous):

\[2.57*2.57-2*2.57*0.57+0.57*0.57*0.57/2.57-0.57\]

OpenStudy (anonymous):

please can you make this simple

ganeshie8 (ganeshie8):

i dont see any pattern. you may let a = 2.57, b = .57 and give a try though..

OpenStudy (anonymous):

kk il giv it a try

OpenStudy (anonymous):

Find the diagonal of a cube whose surface area is 600 cm2

ganeshie8 (ganeshie8):

whats the formula for surface area of cube ?

OpenStudy (anonymous):

6L^2

ganeshie8 (ganeshie8):

6L^2 = 600 L = ?

OpenStudy (anonymous):

l=side